HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

LATEST UPDATES TO THE DIJON CALCULATED SPECTROSCOPIC DATABASE

\underline{C. RICHARD} , V. BOUDON, P. HARDY, Laboratoire Interdisciplinaire Carnot de Bourgogne, UMR 6303 CNRS - Université Bourgogne Franche-Comté, 9 Av. A. Savary, BP 47870, F-21078 Dijon Cedex, France; M. ROTGER, Groupe de

Spectrométrie Moléculaire et Atmosphérique (GSMA), CNRS UMR 7331, Université de Reims Champagne-Ardenne, U.F.R. Sciences Exactes et Naturelles, Moulin de la

Housse B.P. 1039, Reims Cedex 2 51687, France

We report the current status of our calculated spectroscopic relational databases. They contain line lists, as well as cross-section data that are computed by the binned sum of intensities, for specific molecules, that result from recently published spectroscopic analyses. The two original databases, denoted MeCaSDa (CH_{4}) and ECaSDa (C_{2}H_{4}), are regularly improved: the last one should be updated very soon. Then, seven others databases, TFMeCaSDa (CF_{4}), SHeCaSDa (SF_{6}), GeCaSDa (GeH_{4}), RuCaSDa (RuO_{4}), TFSiCasDa (SiF_{4}), UHeCaSDa (UF_{6}) and the last introduced, ChMeCaSDa (CH_{3}Cl), were deployed based upon the same model and can be output into the HITRAN format. These databases are developed in the framework of the international consortium VAMDC (Virtual Atomic and Molecular Data Centre, http://vamdc.org) and are also part of the Dat@OSU project (http://dataosu.obsbesancon.fr).

p-number: p261  Submitted on Wed, 12 Jul 2023 13:45:40 +0200